HUBBLE SPACE TELESCOPE DAILY REPORT #1853 PERIOD COVERED: 0000Z (UTC) 04/09/97 - 0000Z (UTC) 04/10/97 Daily Status Report as of 100/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed NIC/1/2/3 7034 (NICMOS Field Offset Mechanism Test) The NICMOS Near Infrared Cameras (Cameras 1, 2 and 3) were used to = conduct a field offset mechanism test by making observations of P330E. A = 25-point spiral dither pattern is used to generated a 5 x 5 grid of point = source and thermal background observations in each camera. The tests = were conducted as scheduled, and no anomalies were reported. 1.2 Completed WF/PC-2 6731 (Exploring PN Production as a Probe of = Elliptical Galaxy Stellar Populations) The WF/PC-2 was used to make observations of the galaxy NGC-3379 = in order to explore planetary nebulae production as a function of = elliptical galaxy stellar population. In addition to being an excellent = distance indicator, the O III Lambda 5007 A0 planetary nebula luminosity = function is potentially a powerful probe of stellar population within = these galaxies. The observations were executed as scheduled, and no = anomalies were noted. ************************************* Zero gyro safemode occurred at approximately 099/171712Z. See details in Section 2.3. ************************************** The following observations could not be executed due to the safemode = event: 1.3 WF/PC-2 6837 (The Ultraviolet Properties of "Dwarf" Seyfert 1 Nuclei) 1.4 STIS/CCD 7092 (CCD Dark Rate and Read Noise) 1.5 WF/PC-2 6632 (Morphologies of High Redshifted Radio Galaxies) 1.6 NIC 1/2/3 7049 (NICMOS Absolute Photometry Test) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 0 Scheduled Reacquisitions : 8 Successful : 4 All acquisition and re-acquisition "failures" were due to the = safemode entry. 2.2 FHST Updates: Scheduled : 19 Successful : 0 No scheduled updates were performed due to the safemode entry. 2.3 Operations Notes: The battery 6 capacity test was completed at 099/1907Z. During a Type 2 vehicle slew, zero gyro safemode was entered at = approximately 099/171712Z. Telemetry (at 099/155232Z) prior to the slew = indicated that the gyro 4 motor current dropped from 160 mA to 0 mA. = Then, during the slew, anomalous gyro 4 rate output caused gyro disparity = count violations in gyros 3, 5, and 6 which resulted in the safemode = entry. The PSEA was patched for three gyro operations at 099/1921Z, = safemode recovery SPC load was patched for the same reason at 100/0333Z, = and the HST was recovered to software sunpoint at 100/0507Z. Gyro 4 was powered off at 099/2123Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: After a meeting this morning at 8:30L, plans are to go to inertial = hold. Approximately 10-12 hours after this is accomplished, the health = and safety load should commence, and after another 12 hours the intercept = science and SMOV SMS should start. The specifics of these plans are not = currently known. Resumption of normal science and SMOV operations is therefor = expected sometime tomorrow. /DMH